For the given equation, list the intercepts and test for symmetry.
x2+y−121=0



























Answer by stanbon(75887)

For the given equation, list the intercepts and test for symmetry.
x2+y−121=0
-----
f(x) = -x^2-121
f(-x) = -x^2 - 121
-f(-x) = x^2-121
------
Since f(x) = f(-x) you have y-axis symmetry.
----------
Intercepts:
If x = 0, y = -121:: so y-int is (0,-121)
-----
If y = 0, x^2 + -121 ; so x is not a Real Number ; so no x-intercepts.
==================
Cheers,
Stan H.
==================